Below is a graph showing the magnitude of a spring
Answer #1

a) force,f = kx => k = \Deltaf/\Deltax = slope = 5-3/1-0.6 = 2/0.4 = 5 N/m

b) f = mg = kx => x = mg/k =W/k = 28.5/5 = 5.7 m